A 14-year bond with a face value of $1000 is redeemable at twice par and pays coupons semi-annually at j2 = 9.7%. If the yield rate is j2 = 7.6%, find the book value of the bond immediately after the payment of the 11th coupon.
Answer #1

Nper = (14 * 2) - 11= 28 - 11 = 17
Rate = 7.6% / 2 = 3.8%
PMT = 1000 * 9.7% / 2 = 48.5
FV = 1000 * 2 = 2000

Book value of the bond can be calculated by using the following excel formula:
=PV(rate,nper,pmt,fv)
=PV(3.8%,17,-48.5,-2000)
= $1,660.19

Book value of the bond = $1,660.19